OpenStudy (anonymous):

5) WILL GIVE MEDAL AND FAN!! Use the discriminant to determine the number and type of solutions for the following equation. 24x2 – 13 = 0 a. Zero real solutions b. One rational solution c. Two rational solutions d. Two irrational solutions

OpenStudy (anonymous):

The discriminant is b^2 - 4ac. In your problem, a = 24 and b = 0 and c = -13. So, b^2 - 4ac = (0)^2 - 4(24)(-13) = 0 +1248 Since 1248 is greater than 0 but NOT a perfect square, the roots are real, and have irrational solutions.

OpenStudy (anonymous):

ohhh ok so it would be C? @Zackinblack

OpenStudy (anonymous):

Yes sorry was away But yes, it would be C

OpenStudy (anonymous):

You need to first know the definitions of all those bug words. even idk